OpenStudy (anonymous):

Find the number of real number solutions for the equation. x^2-18=0

OpenStudy (jdoe0001):

x^2=18=0 ?

OpenStudy (jdoe0001):

http://www.mathsisfun.com/algebra/fundamental-theorem-algebra.html <--- check the polynomial's degree, THAT MANY solutions

myininaya (myininaya):

There is no solution for x such that x^2 will equal to both 18 and 0 at the same time. Mainly because 18 doesn't equal 0. However I think you mean something like x^2-18=0 or x^2+18=0.

OpenStudy (anonymous):

x^2-18=0 sorry

myininaya (myininaya):

You can actually factor x^2-18 or you can even +18 on both sides and use the square root property to finish solving

myininaya (myininaya):

Do you know how to solve x^2=18?